OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)). Select one: a. -7 b. -3 c. 10 d. 33

OpenStudy (anonymous):

do you know how to find \(g(2)\) ?

OpenStudy (anonymous):

if not, let me know and i will walk you through it

undeadknight26 (undeadknight26):

isn'[t this just: f(x) = x + 8 F(2) = 2 + 8 f(2) = 10?

OpenStudy (anonymous):

no

undeadknight26 (undeadknight26):

Oh sorry then :D

OpenStudy (anonymous):

no i dont

OpenStudy (anonymous):

\[g(x) = x^2 - 6x - 7.\] \[g(2)=2^2-6\times 2-7\]

OpenStudy (anonymous):

i.e. replace \(x\) by \(2\) and see what number you get let me know what it is

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

no i don't think so \[4-12-7\] is not \(1\)

OpenStudy (anonymous):

-15

OpenStudy (anonymous):

ok good then \[f(g(2)=f(-15)\] so your last job is to compute \[f(-15)=-15+8\]

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thank you